# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### 15 Mar 2024 v2

@@ 2 Related Works
State Space Models. Recently, state space sequence models (SSMs) [14, 15], drawing inspiration from classical state-space models [25], have emerged as a promising architecture for sequence modeling. Mamba [12] introduces a selective SSM architecture, integrating time-varying parameters into the SSM framework, and proposes a hardware-aware algorithm to facilitate highly efficient training and inference processes. Some research works leverage SSM in computer vision to process 2D data. The 2D SSM [3] introduces an SSM block at the beginning of each transformer block [51, 8]. This approach aims to achieve efficient parameterization, accelerated computation, and a suitable normalization scheme. SGConvNeXt [29] presents a structured global convolution method inspired by ConvNeXt [32], incorporating multi-scale sub-kernels to achieve both parameterization efficiency and effective long sequence modeling. ConvSSM [46] integrates the tensor modeling principles of ConvLSTM [45] with SSMs, elucidating the utilization of parallel scans in convolutional recurrences. This approach enables subquadratic parallelization and rapid autoregressive generation. Vim [56] introduces a bidirectional SSM block for efficient and versatile visual representation learning, achieving performance comparable to established ViT [8] methods. VMamba [31] introduces a Cross-Scan Module (CSM) designed to traverse the spatial domain and transform any non-causal visual image into ordered patch sequences. This approach achieves linear complexity while preserving global receptive fields. There have also been attempts to utilize SSMs to handle higher-dimensional data. Mamba-ND [28] explores various combinations of SSM and different scan directions within the SSM block to adapt Mamba [12] to higher-dimensional tasks. Recent efforts have sought to replace the traditional transformer-based U-Net within the diffusion denoiser with the SSM block, with the aim of enhancing image generation efficiency. DiffuSSM [52] adeptly manages higher resolutions without relying on global compression, thus maintaining detailed image representation throughout the diffusion process.

### 19 Mar 2024 v3

@@ 2 Related Works
State Space Models. Recently, state space sequence models (SSMs) [14, 15], drawing inspiration from classical state-space models [25], have emerged as a promising architecture for sequence modeling. Mamba [12] introduces a selective SSM architecture, integrating time-varying parameters into the SSM framework, and proposes a hardware-aware algorithm to facilitate highly efficient training and inference processes. Some research works leverage SSM in computer vision to process 2D data. The 2D SSM [3] introduces an SSM block at the beginning of each transformer block [51, 8]. This approach aims to achieve efficient parameterization, accelerated computation, and a suitable normalization scheme. SGConvNeXt [29] presents a structured global convolution method inspired by ConvNeXt [32], incorporating multi-scale sub-kernels to achieve both parameterization efficiency and effective long sequence modeling. ConvSSM [46] integrates the tensor modeling principles of ConvLSTM [45] with SSMs, elucidating the utilization of parallel scans in convolutional recurrences. This approach enables subquadratic parallelization and rapid autoregressive generation. Vim [57] introduces a bidirectional SSM block [52] for efficient and versatile visual representation learning, achieving performance comparable to established ViT [8] methods. VMamba [31] introduces a Cross-Scan Module (CSM) designed to traverse the spatial domain and transform any non-causal visual image into ordered patch sequences. This approach achieves linear complexity while preserving global receptive fields. There have also been attempts to utilize SSMs to handle higher-dimensional data. Mamba-ND [28] explores various combinations of SSM and different scan directions within the SSM block to adapt Mamba [12] to higher-dimensional tasks. Recent efforts have sought to replace the traditional transformer-based U-Net within the diffusion denoiser with the SSM block, with the aim of enhancing image generation efficiency. DiffuSSM [53] adeptly manages higher resolutions without relying on global compression, thus maintaining detailed image representation throughout the diffusion process.
@@ 5 Discussion and Conclusion
Motion Mamba Supplementary
(a) Text-Motion Correspondence User Study
(b) Quality User Study
Figure 5: User Study in two aspects including text-motion correspondence and quality, we compare Motion Mamba(MM) with previous methods including MDM [49] , T2M [17] , MLD [6] and ground truth.
6 Implementation Details
Motion Mamba operates within the latent spaces, leveraging the capabilities of the Motion Variational AutoEncoder (VAE) $\mathcal{V}=\{\mathcal{E},\mathcal{D}\}$, as proposed in the seminal work by Chen et al. [6]. For the configuration of the Motion Mamba denoiser $\epsilon_{\theta}$, we have opted for an architecture comprising 11 layers ($N=11$), with the latent dimensionality set to $z\in\mathbb{R}^{2,d}$. The Hierarchical Temporal Mamab (HTM) modules are arranged in a scan pattern of $\{S^{2N_{n}-1},\ldots,S^{1}\}$, while the Bidirectional Spatial (BSH) modules incorporate a block-level bidirectional scan policy. Additionally, we utilize a pretrained CLIP-VIT-L-14 model in a frozen state to derive text embeddings $\tau_{\theta}^{w}(w^{1:N})\in\mathbb{R}^{1\times d}$.
All models under the Motion Mamba framework are meticulously trained using the AdamW Optimizer, with the learning rate steadfastly maintained at $10^{-4}$. We have standardized our global batch size at 512, which is judiciously distributed across 4 GPUs to facilitate data-parallel training. The training regime is extended over 2,000 epochs to ensure convergence to an optimal set of parameters. For the diffusion sampling process, we maintain the number of steps at 1,000 and 50 during the training and inference phases, respectively. The entire training procedure is executed on a single-node GPU server, outfitted with 4 NVIDIA A100 GPUs, spanning approximately 4 hours. Inference speed evaluations of our Motion Mamba models are conducted on a single NVIDIA V100 GPU for fair comparison, while module development and additional inference tasks are performed on a single NVIDIA GeForce RTX 3090/4090 GPU.
7 User Study
In this work, we undertake a comprehensive evaluation of Motion Mamba’s performance, encompassing both qualitative analyses across various datasets and a user study to assess its real-world applicability. A diverse collection of 20 motion sequence sets, prompted randomly and extracted from the HumanML3D [17] test set, were generated utilizing three distinct methodologies—MDM [49], T2M [17], MLD [6]—alongside Motion Mamba and a baseline of ground truth motions. Subsequently, 50 participants were randomly selected to evaluate the motion sequences generated by these methods.
The user study was administered through a Google Forms interface, as depicted in Fig. 8, ensuring that motion sequences were presented anonymously without revealing their generative model origins. Our analysis focused on two critical dimensions: the fidelity of text-to-motion correspondence and the overall quality of the generated motions.
Empirical results, illustrated in Fig. 4(a) and Fig. 4(b), unequivocally demonstrate Motion Mamba’s superior performance relative to the benchmark methods in terms of both text-motion alignment and motion quality. Specifically, Motion Mamba achieved significant margins over MDM [49], T2M [17], and MLD [6] by 79%, 74%, and 62% in text-motion correspondence, respectively, as highlighted in Fig. 4(a). When juxtaposed with ground truth data—meticulously captured with state-of-the-art, noise-free devices—Motion Mamba’s generated sequences exhibited a remarkably close adherence to the intended text descriptions, underscoring its proficiency in aligning textual prompts with motion sequences.
Further reinforcing these findings, Motion Mamba’s generated motions were also found to surpass the aforementioned methods by substantial margins of 70%, 67%, and 59%, respectively, in terms of quality, as reported in Fig. 4(b) . This underscores Motion Mamba’s ability to not only closely match the text-motion correspondence of high-fidelity ground truth data but also to produce high-quality motion sequences that resonate well with real user experiences.
8 Visualization
Our study delves into the visualization of motion generation by capturing intricate motion sequences, utilizing prompts and their variations derived from HumanML3D [17]. We meticulously compare our proposed Motion Mamba methodology with established state-of-the-art techniques, namely MotionDiffuse [54], MDM [49], and MLD [6]. Presenting three distinct motion sequences, we meticulously analyze and visualize each, offering a comprehensive assessment of our approach’s efficacy.
Figure 6: We compared the proposed Motion Mamba with well-established state-of-the-art methods such as MotionDiffuse [54], MDM [49], and MLD [6]. We presented three distinct motion prompts and visualized them in the form of motion sequence. The results demonstrated our superior performance compared to existing methods.
Figure 7: We have included extra examples to showcase the proposed Motion Mamba model. These examples feature randomly selected prompts sourced from HumanML3D [17], providing additional visualizations of the model’s capabilities.
Figure 8: This figure presents the User Interface (UI) deployed for our User Study, wherein participants are presented with two videos, labeled as Video A and Video B, respectively. These videos are selected randomly from a pool consisting of outputs generated by three distinct methods, in addition to the Ground Truth (GT) for comparison. Participants are posed with two types of evaluative questions to gauge the effectiveness of the generated motions. The first question, "Which of the two motions is more realistic?", aims to assess the overall quality and realism of the motion capture. The second question, "Which of the two motions corresponds more accurately to the text prompt?", is designed to evaluate the congruence between the generated motion and the provided text prompt. This dual-question approach facilitates a comprehensive assessment of both the quality of the motion generation and its fidelity to the specified text prompts.